# From Stimuli to Minds: Enhancing Psychological Reasoning in LLMs via Bilateral Reinforcement Learning

### v1

@@ Abstract
Large Language Models show promise in emotion understanding, social reasoning, and empathy, yet they struggle with psychologically grounded tasks that require inferring implicit mental states in context-rich, ambiguous settings. These limitations arise from the absence of theory-aligned supervision and the difficulty of capturing nuanced mental processes in real-world narratives. To address this gap, we leverage expert-labeled, psychologically rich scenarios and propose a trajectory-aware reinforcement learning framework that explicitly imitates expert psychological thought patterns. By integrating real-world stimuli with structured reasoning guidance, our approach enables compact models to internalize social-cognitive principles, perform nuanced psychological inference, and support continual self-improvement. Comprehensive experiments across multiple benchmarks further demonstrate that our models achieve expert-level interpretive capabilities, exhibiting strong out-of-distribution generalization and robust continual learning across diverse, challenging, and psychologically grounded tasks.
@@ 1 Introduction
Large Language Models (LLMs) show strong generalization across diverse language tasks (Wu 2025; Zhang et al. 2025; Feng et al. 2025). Their emerging potential in psychological domains, such as emotion understanding (Kovacevic et al. 2024), social reasoning (Leng and Yuan 2023), and empathy recognition (Sorin et al. 2024) has drawn growing research interest. These tasks require not only linguistic comprehension but also nuanced inference of implicit mental states and emotional cues, often without explicit supervision or clearly defined ground truths. Unlike conventional language tasks, they involve rich psychological stimuli embedded in ambiguous, socially grounded scenarios shaped by diverse cultural norms, interpersonal dynamics, and lived experiences. Although LLMs show partial sensitivity to such cues, performance on psychologically grounded tasks remains limited from human-level competence (Ke et al. 2025).
Figure 1: A sample ToMbench question presents two psychological stimuli—one focused on direct emotion prediction, the other requiring nuanced reasoning about mental states.
Figure 2: The figure summarizes key psychological stimuli across our datasets—emotional variables, narrative transformations, and collective factors—comprising 35,084, 24,831, and 36,890 QA instances, each offering insights into human narratives.
@@ 2.1 LLMs in Psychological Tasks
LLMs are increasingly evaluated on psychological tasks involving theory of mind and moral reasoning. Benchmarks show that even top models such as GPT-4o underperform humans in belief reasoning, indicating reliance on superficial cues (Xiao et al. 2025; Shapira et al. 2024). Datasets such as SimpleToM, CogBench, and SocialIQa extend evaluations to emotion and social inference (Gu et al. 2024; Coda-Forno et al. 2024; Sap et al. 2019b). In moral reasoning, LLMs may align with or even surpass humans in perceived ethical competence (Dillion et al. 2025; Huang et al. 2023), though often via pattern matching. Empathy studies find LLMs capable of emotionally appropriate responses, occasionally preferred over human ones (Ayers et al. 2023; Sorin et al. 2024), albeit with limited contextual nuance (Yang, Ye, and Du 2024). Open models still lag behind proprietary ones (Li et al. 2024), highlighting the need for theory-informed benchmarks and training methods (Xie et al. 2024; Qiu et al. 2024).
@@ 2.2 CoT and Reinforcement Learning
Chain-of-Thought (CoT) prompting improves deliberate reasoning by encouraging stepwise thinking (Wei et al. 2022; Kojima et al. 2022), and is often combined with RL to align outputs with human preferences (Ouyang et al. 2022). Process supervision and gradient-level feedback, such as in DeepSeek-R1, further enhance intermediate reasoning quality (Liu and et al. 2025). Frameworks like Reflexion and Tree of Thoughts introduce self-evaluation and planning to strengthen coherence (Shinn et al. 2023; Yao et al. 2023). Despite progress, challenges remain in balancing fluency with rigor and designing scalable feedback for ambiguous reasoning. CoT and RL provide a strong basis for aligning LLMs in psychological reasoning tasks and broader real-world applications.
@@ 3.2 Bilateral Reinforcement Learning
where $y^{\text{pred}}_{b,i}$ $y^{\text{gold}}_{b,i}$ are token sets after normalization, including lowercase conversion, punctuation and space cleanup.
@@
[…]
@@ 3.3 Continuous Learning
To enable self-evaluation and self-improvement under low-resource settings, we propose a simple framework for continuous learning. The model not only generates predictions but also assesses their quality and improves over time—producing more user-aligned and psychologically appropriate responses. We define the learning criterion as:
[…]
@@ 4.2 RQ1: Effect of Human-Labeled Data
One of the central hypotheses in our work is that human-labeled data, particularly when annotated by trained psychology students, can inject high-quality domain knowledge into compact models during SFT. We evaluate this hypothesis by comparing models trained on our StimuliQA dataset against those trained on synthetic datasets generated by two powerful large models: Llama 3.3-70B and Mistral 8$\times$7B. All datasets were used to train Qwen2.5-Instruct models of three different sizes (0.5B, 1.5B, 3B) via SFT, and evaluated on five diverse OOD benchmarks: ToMbench, SimpleToM, SocialIQa, CosmosQA, and BIG-Bench Hard. Where we have the final results displayed in Table 1 and Figure 4. This comparison reveals performance gaps and tests the transferability of psychological reasoning across model sizes and tasks. These benchmarks span theory of mind, social commonsense, and logical reasoning, allowing for a comprehensive assessment of domain generalization and model capabilities.
As shown in Table 1, models trained on StimuliQA consistently outperform those trained on synthetic data across all model sizes and benchmarks, with especially strong gains in the 3B setting. For instance, on SimpleToM, Qwen2.5-3B achieves 37.62 F1 and 56.44% accuracy, far exceeding Llama 3.3 (18.48/26.16%) and Mistral 8$\times$7B (35.02/33.33%). These substantial margins further reinforce our claim that: human-labeled data conveys more nuanced and structured psychological knowledge than synthetic outputs. The consistent OOD improvements indicate this knowledge is indeed generalizable, highlighting the professional quality and domain relevance of our dataset. Annotations by trained psychology students provide valuable domain expertise essential for effective model training. When used in GRPO training (Figure 4), human-labeled data still offers an advantage.
@@ Appendix B Experimental Details
We build our experimental datasets starting from StimuliQA, which contains 3,280 expert-annotated narratives covering 54 psychological variables. To increase topic diversity and mitigate repetition, we manually extract 10–50 candidate topics per variable based on sample size, and then expand them into up to 200 diverse topics each using in-context learning (ICL) prompts with LLaMA 3.3 Instruct 70B. Following the same pipeline, we construct two external comparison datasets: MistralQA (from Mistral 8$\times$7B) and LlamaQA (from LLaMA 3.3 Instruct 70B). These datasets collectively provide a broad coverage of psychological reasoning scenarios for pre-training and evaluation.
@@ Appendix C Dataset and Code
Please check our anonymous repo via the link https://anonymous.4open.science/r/Stimuli2Minds-E37B for our available datasets and code. Our experiment results will be further listed in this given repository.
@@ C.3 Model reasoning sample
Mrs. Li should be proud of her son becoming the football team captain, but why does she actually seem worried?
@@
Reasoning: Mrs. Li might be worried for a few reasons. First, she may be concerned about the pressure Xiao Qiang faces as a captain, including the expectations from his teammates, coaches, and school administration. Second, she might be worried about his safety and the risks associated with playing football, especially if she is concerned about injuries. Additionally, she might be worried about his academic performance, as balancing sports and studies can be challenging. Lastly, she might be concerned about his well-being and the emotional toll of high-pressure situations.
@@
Bilateral Answer: Because she is worried about his safety and the pressure of the role.
Table 12: Comparison between the Base model, Bilateral model, and ground truth on a ToMbench sample (Index2).
As shown in Table 12, both the Base and Bilateral interpretations attempt to explain Mrs. Li’s worry despite her pride in Xiao Qiang’s achievement. The Base model produces a long, enumerative reasoning sequence, speculating on multiple possible concerns—academic performance, physical safety, emotional stress, and social expectations. While this reasoning is broadly plausible, it reflects a diffuse cognitive pattern in psychological terms, where the model distributes attention across many secondary emotional states instead of isolating the dominant source of worry. This over-generation risks introducing irrelevant inferences that dilute alignment with the ground truth.
In contrast, the Bilateral Interpretation adopts a focused and dual-channel reasoning strategy aligned with social-cognitive psychology. It highlights the two most salient drivers of parental worry: role pressure and safety concerns. These correspond to the primary appraisal dimensions in Lazarus’ cognitive-motivational-relational theory of emotion, where parents appraise both the demands of the new social role and the potential risks it entails. By selectively attending to the most psychologically plausible causes of worry, the Bilateral model exhibits higher signal-to-noise ratio in emotional inference and aligns more closely with the ground truth reasoning process.
@@ Appendix D Limitation
While our current approach leverages length‑based reasoning incentives to guide LLMs toward more structured and cognitively transparent reasoning in psychological tasks, it remains limited in granularity. Relying solely on length encourages step‑by‑step reasoning but may not fully capture the qualitative dimensions of psychological inference, such as the emotional salience or contextual accuracy of mental‑state predictions. A promising future direction is to incorporate fine‑grained reward signals—for example, rewarding accurate emotional appraisal, causal consistency, or alignment with psychological theory—so that the LLM receives multi‑dimensional reinforcement rather than a single length‑based signal. This would enable more precise and nuanced guidance in modeling human‑like mental‑state reasoning.

### v4

@@ Abstract
Large Language Models show promise in emotion understanding, social reasoning, and empathy, yet they struggle with psychologically grounded tasks that require inferring implicit mental states in context-rich, ambiguous settings. These limitations arise from the absence of theory-aligned supervision and the difficulty of capturing nuanced mental processes in real-world narratives. To address this gap, we leverage expert-labeled, psychologically rich scenarios and propose a trajectory-aware reinforcement learning framework that explicitly imitates expert psychological thought patterns. By integrating real-world stimuli with structured reasoning guidance, our approach enables compact models to internalize social-cognitive principles, perform nuanced psychological inference, and support continual self-improvement. Experiments across multiple benchmarks demonstrate that our models achieve expert-level interpretive capabilities, exhibiting strong out-of-distribution generalization across diverse psychologically tasks. Our code is publicly available at https://github.com/Githubuseryf/Stimuli2Minds.
@@ 1 Introduction
Large Language Models (LLMs) show strong generalization across diverse language tasks (Wu 2025; Zhang et al. 2025; Feng et al. 2025b). Their emerging potential in psychological domains, such as emotion understanding (Kovacevic et al. 2024), social reasoning (Leng and Yuan 2023), and empathy recognition (Sorin et al. 2024) has drawn growing research interest. These tasks require not only linguistic comprehension but also nuanced inference of implicit mental states and emotional cues, often without explicit supervision or clearly defined ground truths. Unlike conventional language tasks, they involve rich psychological stimuli embedded in ambiguous, socially grounded scenarios shaped by diverse cultural norms, interpersonal dynamics, and lived experiences. Although LLMs show partial sensitivity to such cues, performance on psychologically grounded tasks remains limited from human-level competence (Ke et al. 2025).
Figure 1: A sample ToMbench question presents two types of psychological stimuli pairs for demonstration.
Figure 2: The figure summarizes key psychological parameters across our datasets: emotional variables, narrative transformations, and collective factors, comprising 35,084, 24,831, and 36,890 QA instances, each offering insights into human narratives.
@@ 2.1 LLMs in Psychological Tasks
LLMs are increasingly evaluated on psychological tasks involving theory of mind and moral reasoning. Benchmarks show that even top models such as GPT-4o underperform humans in belief reasoning, indicating reliance on superficial cues (Xiao et al. 2025; Shapira et al. 2024). Datasets such as SimpleToM, CogBench, and SocialIQa extend evaluations to emotion and social inference (Gu et al. 2024; Coda-Forno et al. 2024; Sap et al. 2019b). In moral reasoning, LLMs may align with or even surpass humans in perceived ethical competence (Liu et al. 2025; Huang et al. 2023), though often via pattern matching. Empathy studies find LLMs capable of emotionally appropriate responses, occasionally preferred over human ones (Ayers et al. 2023; Sorin et al. 2024), albeit with limited contextual nuance (Yang, Ye, and Du 2024). Open models still lag behind proprietary ones (Li et al. 2024), highlighting the need for theory-informed benchmarks and training methods (Xie et al. 2024; Qiu et al. 2024).
@@ 2.2 CoT and Reinforcement Learning
Chain-of-Thought (CoT) prompting improves deliberate reasoning by encouraging stepwise thinking (Luo et al. 2025a; Shen et al. 2025), and is often combined with RL to align outputs with human preferences (Ouyang et al. 2022). Process supervision and gradient-level feedback, such as in DeepSeek-R1, further enhance intermediate reasoning quality (Luo et al. 2025c). Frameworks like Reflexion and Tree of Thoughts introduce self-evaluation and planning to strengthen coherence (Luo et al. 2025b; Feng et al. 2025a). Despite progress, challenges remain in balancing fluency with rigor and designing scalable feedback for ambiguous reasoning. CoT and RL provide a strong basis for aligning LLMs in psychological reasoning tasks and broader real-world applications.
@@ 3.2 Bilateral Reinforcement Learning
where $y^{\text{pred}}_{b,i}$ & $y^{\text{gold}}_{b,i}$ are token sets after normalization, including lowercase conversion, punctuation and space cleanup.
@@
To encourage informative yet concise reasoning, we introduce a bilateral reward term that jointly considers relative reasoning length and answer quality. Let $L_{b,i}=|o_{b,i}|$ denote the output length for group $(b,i)$, and define the batch-level averages as:
@@ 3.3 Continuous Learning
To enable self-evaluation and self-improvement under low-resource settings, we propose a simple framework for continuous learning. The model not only generates predictions but also assesses their quality and improves over time, producing more user-aligned and psychologically appropriate responses. We define the learning criterion as:
$\text{self\_train}(x)\iff\text{valid}(x)\land\text{confidence}(x)>\tau,$ (10)
@@ 4.2 RQ1: Effect of Human-Labeled Data
[…]
@@ Appendix B Experimental Details
[…]
@@ Appendix C Dataset and Code
Our experiment results will be further listed in our repository.
@@ C.3 Model reasoning sample
Mrs. Li’s son, Xiao Qiang, is now in college. He tells her that he has become the captain of the school football team, a notable achievement since the team is highly popular on campus.
@@
Although Mrs. Li should feel proud of her son becoming the football team captain, why does she seem worried?
@@
Reasoning: Mrs. Li may be worried for several reasons. She might be concerned about the pressure Xiao Qiang faces as a team captain, including high expectations from teammates and coaches. She may also worry about his physical safety and the risk of injuries, as well as the challenge of balancing football responsibilities with his academic workload. Additionally, she could be anxious about the emotional toll that high-pressure situations might bring.
@@
Reasoning: Mrs. Li is mainly worried that her son will struggle to balance his studies with his demanding role as team captain, and that the added pressure of leadership may affect his well-being.
@@
Bilateral Answer: Because she is worried about the pressure of leadership and its impact on his studies.
Table 12: Comparison of Base model, Bilateral model, and ground truth reasoning on a ToMbench sample (Index 2).
As shown in Table 12, both the Base and Bilateral models correctly identify that Mrs. Li’s worry stems from the tension between pride and concern. The Base model produces a lengthy, enumerative reasoning chain, speculating about multiple secondary concerns such as injuries, social expectations, and emotional stress. While plausible, this diffuse reasoning pattern risks diluting alignment with the ground truth by introducing peripheral inferences.
By contrast, the Bilateral model exhibits a focused reasoning strategy, isolating the dominant source of worry: the pressure of leadership and its impact on academic balance. This reasoning reflects a higher signal-to-noise ratio in emotional inference and aligns more closely with the ground truth, illustrating how bilateral reinforcement guides the model toward psychologically salient appraisals rather than exhaustive speculation.
@@ Appendix D Limitation
While our current approach leverages length-based reasoning incentives to guide LLMs toward more structured and cognitively transparent reasoning in psychological tasks, it remains limited in granularity. Relying solely on length encourages step-by-step reasoning but may not fully capture the qualitative dimensions of psychological inference, such as emotional salience or contextual accuracy in mental-state predictions. A promising direction is to incorporate fine-grained reward signals—for example, rewarding accurate emotional appraisal, causal consistency, or alignment with psychological theory—so that the LLM receives multi-dimensional reinforcement rather than a single length-based signal.
Integrating such rewards could help the model balance interpretability with empathy, learning not only to elaborate its reasoning but also to understand why certain inferences are more plausible or emotionally resonant. Exploring hierarchical reward structures that jointly optimize structural reasoning and affective understanding may lead to a more holistic, human-aligned reasoning process.